# NN-ResDMD: Learning Koopman Representations for Complex Dynamics with Spectral Residuals

### v1

@@ A.4 Highlights of NN-ResDMD compared with typical existing neural network-based Koopman framework
Our NN-ResDMD method takes a fundamentally different approach from existing deep learning methods by building upon the residual-based framework of ResDMD rather than the different Koopman-approximating loss functions following the variational principles of VAMPnets [39, 40, 29] or the deep autoencoder structure in [41]. By incorporating spectral residual measures into deep learning and introducing a structured representation that captures dependencies among eigenvalues, we achieve more compact and interpretable models for nonlinear systems with continuous spectra. This approach enables us to directly minimize Koopman spectral approximation errors while avoiding the high-dimensional representations or point-spectrum limitations of previous methods.
@@
If we take the VAMP framework as an example, here are the connections and differences. The proposed loss function and the VAMP score share the goal of optimizing approximations of the Koopman operator’s spectral properties, establishing a connection in their ultimate purpose. However, although they both depend on the covariance matrices (in our manuscript Equation 3.2), their methodologies differ significantly. Our residual-based method directly minimizes the spectral approximation error of the Koopman operator and accommodates both point and continuous spectra, while the VAMP score follows a variational framework, maximizing the sum of singular values to approximate the point spectrum, primarily for stochastic systems (though see an exception in TianWu+2021+635+659 [39]). Moreover, while VAMP is specifically designed for Markov processes and requires the Koopman operator to be Hilbert-Schmidt, our approach focuses on deterministic systems and enables a more comprehensive spectral analysis that incorporates continuous spectra. This distinction in scope and methodology highlights how the two frameworks complement each other in addressing different aspects of spectral estimation.
A.5 Discussion of computation costs
@@
Despite the various advantages of the NN-ResDMD framework, one significant limitation is its higher computational cost compared to the original ResDMD and other classical methods.
@@
Theoretical Perspective: The NN-ResDMD algorithm’s computational demands stem primarily from its iterative optimization process. Each iteration involves a gradient descent update with complexity scaling linearly with both system dimensionality and neural network parameters. Though individual gradient steps are computationally lightweight for standard network architectures, the algorithm’s efficiency issue lies in its repeated least-squares optimizations. Below we provide a detailed comparison between computation costs of several methods we used in this paper.
@@
Comparison over methods: The computational costs of EDMD, ResDMD, EDMD-DL, Hankel-DMD, and NN-ResDMD vary significantly based on their core computational steps and specific configurations. For a dataset with $m=10^{5}$ data points and $N_{K}=300$ dictionary functions (for EDMD-based methods), the theoretical complexity and runtime differ across methods. EDMD involves least squares and eigenvalue decomposition, with a complexity of $O(N_{K}^{2}m+N_{K}^{3})$, making it the fastest method, and typically requiring only seconds to minutes for computation. ResDMD extends EDMD by adding residual evaluation and pseudospectrum computation. The residual evaluation introduces an additional $O(N_{K}^{3})$, and pseudospectrum computation across $n_{z}$ grid points incurs $O(n_{z}N_{K}^{3})$, resulting in a total complexity of $O(N_{K}^{2}m+n_{z}N_{K}^{3})$. This leads to runtimes ranging from minutes to hours, depending on the resolution of the pseudospectrum grid. EDMD-DL incorporates dictionary learning through stochastic gradient descent (SGD), where each iteration involves matrix construction ($O(N_{K}^{2}m)$), Koopman matrix computation ($O(N_{K}^{3})$), and neural network forward/backward propagation ($O(d|H|)$, with $d|H|$ representing the total network parameter size). With $k$ SGD iterations, the total complexity becomes $O(k(N_{K}^{2}m+N_{K}^{3}+d|H|))$, leading to runtimes also in the range of minutes to hours depending on $k$. NN-ResDMD, which builds on EDMD-DL, shares the same complexity, $O(k(N_{K}^{2}m+N_{K}^{3}+d|H|))$, but includes the explicit use of Koopman matrix eigenvectors and optional pseudospectrum computation, making its runtime slightly longer than EDMD-DL for high-resolution spectral analysis. Hankel-DMD, using a time delay embedding dimension $T$, constructs a Hankel matrix ($O(Tm)$), performs singular value decomposition (SVD) ($O(T^{2}m)$), and computes the eigenvalues of a reduced $T\times T$ matrix ($O(T^{3})$). The total complexity is $O(Tm+T^{2}m+T^{3})$, and the runtime is heavily influenced by $T$, typically ranging from minutes to hours. While EDMD is computationally the most efficient, ResDMD and Hankel-DMD provide higher precision and robustness in spectral analysis at the expense of increased runtime, and EDMD-DL and NN-ResDMD offer flexibility and accuracy through dictionary learning, with additional SGD iterations and optional pseudospectrum computation contributing to their computational burden.
@@
Empirical Perspective: In our experiments, without computing the pseudospectrum, the computational cost of ResDMD typically ranges from seconds to minutes. NN-ResDMD, on the other hand, can require tens of minutes to several hours, depending on factors such as data dimensionality, the number of snapshots, hidden layer configurations, dictionary sizes, and training convergence criteria.
@@
Trade-off Between Cost and Accuracy: While NN-ResDMD’s additional computational steps introduce higher costs, they enhance the accuracy and robustness of Koopman eigenpair estimation by allowing automatic dictionary learning and minimizing spurious spectral components. This trade-off makes NN-ResDMD particularly valuable in applications where precision is critical. However, its computational demands render it less suitable for real-time or online Koopman model learning tasks.
A.6 Hankel-DMD
A.6.1 Justification of Using Hankel-DMD as comparison in all experiments
@@
As results, although its performance rivals NN-ResDMD in the simple pendulum system by showing eigenvalues points near the unit circle and containing some polluted eigenvalues, which are close to the ground truth unit circle, we would like to emphasize that it capture the point spectrum and miss the full spectral information. When it comes to high-dimensional systems, it fails to capture key dynamics in higher-dimensional systems, as seen in the later experiment (Section 4.2 and Section 4.3).
A.6.2 Application in Turbulence
Here we present the Koopman modes computed by Hankel-DMD for comparison with the NN-ResDMD results. As shown in Figure 7, despite having small residuals, these modes fail to clearly capture the fundamental pressure field structure that was successfully identified by NN-ResDMD’s first Koopman mode (see Figure 5). This comparison demonstrates the superior ability of NN-ResDMD to extract physically meaningful patterns from complex fluid systems.
Figure 7: The plots illustrate turbulence detection using the four Koopman modes computed by Hankel-DMD, which are ranked with their corresponding residuals from the smallest.
A.7 Practical details for neural data analysis
A.7.1 Dataset Details and Experimental Setup
@@
A.7.2 Task Definition and Rationale
In contrast to the competition’s prediction objective, our study focuses on the task of state partitioning of neural signals. While prediction remains feasible, we aim to demonstrate that state partitioning is sufficient to highlight the superiority of NN-ResDMD over a series of other methods in uncovering the latent dynamics of the system. Specifically, in each experiment, a set of six video stimuli was repeatedly presented to each mouse, creating ideal conditions for defining brain states. The recording setup remained consistent for each mouse, ensuring that the neural activities could be interpreted as originating from the same dynamical system, with the primary variable being the input stimulus.
@@ A.7.2 Task Definition and Rationale
A.7.3 Dataset Structure and Dimensionality
@@
A.7.4 Implementations of NN-ResDMD and other classical methods
We compare here four methods: the proposed NN-ResDMD and three classical Koopman decomposition methods for high-dimensional systems: the Hankel-DMD, the EDMD with RBF basis, and the Kernel ResDMD. We applied them to the 5 datasets, although with slightly different implementations and different dimensions of approximated Koopman invariance subspace.
For NN-ResDMD, we train the dictionaries with all the snapshots recorded in each mouse such that the total snapshot number is the product of the snapshot number in one trial and the number of all trials. This is to avoid overfitting with the small snapshot numbers within a trial. The high-dimensional data is first reduced to 300 dimensions with Singular Value Decomposition. The dimension of the Koopman subspace is chosen to be 601, consisting of 300 trained bases and 301 pre-chosen ones (constant and the first-degree polynomials of the SVD-ed 300 dimensions). The first 501 eigenfunctions sorted by the modulus of eigenvalues are selected to avoid spurious eigenvalues estimation due to noise. One can find the decomposed eigenfunctions in Figure 6A(top), with a marker of the ground truth state separations based on stimulus identity.
For Hankel-DMD, the Koopman eigenfunctions were approximated using the eigenvectors of the Hankel matrix. Specifically, the Hankel matrix was formed as in Equation 53 from arbabi2017ergodic [25], using all the observables from one trial of each mouse with a delay of 50. Consequently, the snapshot size became 249 times the observable number, and the resulting number of eigenfunctions was 50, each with a length of 50. The Hankel-DMD eigenfunctions for each trial of data are shown in Figure 6A (bottom), alongside the ground truth trial identities for comparison.
@@ A.7.4 Implementations of NN-ResDMD and other classical methods
The Koopman eigenfunctions from both NN-ResDMD and other methods represent dynamical features corresponding to one of the six video stimuli. To evaluate how well the eigenfunctions capture the latent dynamics, we assess the similarity of the features for trials with the same stimulus and their dissimilarity from those corresponding to different stimuli. Effectively, this makes the problem a clustering task, where the separability of the Koopman eigenfunctions reflects how well they capture the key dynamic components related to the stimuli.
A.8 Choice justification of dictionary sizes
@@
    For NN-ResDMD, we selected 300 trained basis functions and 300 first-order monomial basis functions as the dictionary for the 300 reduced observables. This choice ensures the dictionary is rich enough to span the Koopman invariant subspace. Hence, the size of the trained dictionary was set to be at least equal to the original observable size. Then based on the rank of estimated Koopman eigenvalues, we select the dominant 501 eigenfunctions to avoid the eigenfunctions with zero eigenvalues.
@@
    For Hankel DMD, the number of delays (as dictionary size/number of eigenfunctions) is first constrained by the temporal sample size (i.e., snapshot size) because it cannot exceed the maximum snapshot size. Therefore, it is impossible to choose the same dictionary size as the NN-ResDMD example. Choosing the delay too small will result in an insufficient dictionary size to span the Koopman invariant subspace, and too large will reduce the actual snapshot size to estimate the covariance matrices in the estimation of the Koopman matrix. Therefore, we chose a compromise delay number of 50 that satisfies both needs.
@@
    For RBF basis, in principle, we can use the same dictionary size. However, our previous experience with a similar dataset and the results of using the RBF basis for the EDMD method all suggest that the performance will be better with more dictionary functions. Therefore, we chose 1000 RBF basis and the original 300 first-order monomial basis as a better condition compared to the same dictionary size with NN-ResDMD.
@@
    For Kernel ResDMD, the dictionary size is theoretically determined to be the number of snapshots. Therefore, we cannot make the dictionary size consistent with the NN-ResDMD example.
@@
A.8.1 Visualization and Clustering Performance

### v3

@@ A.5 Highlights of ResKoopNet compared with typical existing neural network-based Koopman framework
Our ResKoopNet method takes a fundamentally different approach from existing deep learning methods by building upon the residual-based framework of ResDMD rather than the different Koopman-approximating loss functions following the variational principles of VAMPnets [37, 45, 25] or the deep autoencoder structure in [24]. By incorporating spectral residual measures into deep learning and introducing a structured representation that captures dependencies among eigenvalues, we achieve more compact and interpretable models for nonlinear systems with continuous spectra. This approach enables us to directly minimize Koopman spectral approximation errors while avoiding the high-dimensional representations or point-spectrum limitations of previous methods.
@@
If we take the VAMP framework as an example, here are the connections and differences. The proposed loss function and the VAMP score share the goal of optimizing approximations of the Koopman operator’s spectral properties, establishing a connection in their ultimate purpose. However, although they both depend on the covariance matrices (in our manuscript Equation 3.2), their methodologies differ significantly. Our residual-based method directly minimizes the spectral approximation error of the Koopman operator and accommodates both point and continuous spectra, while the VAMP score follows a variational framework, maximizing the sum of singular values to approximate the point spectrum, primarily for stochastic systems (though see an exception in Tian and Wu [37]). Moreover, while VAMP is specifically designed for Markov processes and requires the Koopman operator to be Hilbert-Schmidt, our approach focuses on deterministic systems and enables a more comprehensive spectral analysis that incorporates continuous spectra. This distinction in scope and methodology highlights how the two frameworks complement each other in addressing different aspects of spectral estimation.
A.6 Discussion of computation costs
@@
Despite the various advantages of the ResKoopNet framework, one significant limitation is its higher computational cost compared to the original ResDMD and other classical methods.
@@
Theoretical Perspective: The ResKoopNet algorithm’s computational demands stem primarily from its iterative optimization process. Each iteration involves a gradient descent update with complexity scaling linearly with both system dimensionality and neural network parameters. Though individual gradient steps are computationally lightweight for standard network architectures, the algorithm’s efficiency issue lies in its repeated least-squares optimizations. Below we provide a detailed comparison between computation costs of several methods we used in this paper.
@@
Comparison over methods: The computational costs of EDMD, ResDMD, EDMD-DL, Hankel-DMD, and ResKoopNet vary significantly based on their core computational steps and specific configurations. For a dataset with $m=10^{5}$ data points and $N_{K}=300$ dictionary functions (for EDMD-based methods), the theoretical complexity and runtime differ across methods. EDMD involves least squares and eigenvalue decomposition, with a complexity of $O(N_{K}^{2}m+N_{K}^{3})$, making it the fastest method, and typically requiring only seconds to minutes for computation. ResDMD extends EDMD by adding residual evaluation and pseudospectrum computation. The residual evaluation introduces an additional $O(N_{K}^{3})$, and pseudospectrum computation across $n_{z}$ grid points incurs $O(n_{z}N_{K}^{3})$, resulting in a total complexity of $O(N_{K}^{2}m+n_{z}N_{K}^{3})$. This leads to runtimes ranging from minutes to hours, depending on the resolution of the pseudospectrum grid. EDMD-DL incorporates dictionary learning through stochastic gradient descent (SGD), where each iteration involves matrix construction ($O(N_{K}^{2}m)$), Koopman matrix computation ($O(N_{K}^{3})$), and neural network forward/backward propagation ($O(|\theta|)$, with $|\theta|$ representing the total network parameter size). With $k$ SGD iterations, the total complexity becomes $O(k(N_{K}^{2}m+N_{K}^{3}+|\theta|))$, leading to runtimes also in the range of minutes to hours depending on $k$. ResKoopNet, which builds on EDMD-DL, shares the same complexity, $O(k(N_{K}^{2}m+N_{K}^{3}+|\theta|))$, but includes the explicit use of Koopman matrix eigenvectors and optional pseudospectrum computation, making its runtime slightly longer than EDMD-DL for high-resolution spectral analysis. Hankel-DMD, using a time delay embedding dimension $T$, constructs a Hankel matrix ($O(Tm)$), performs singular value decomposition (SVD) ($O(T^{2}m)$), and computes the eigenvalues of a reduced $T\times T$ matrix ($O(T^{3})$). The total complexity is $O(Tm+T^{2}m+T^{3})$, and the runtime is heavily influenced by $T$, typically ranging from minutes to hours. While EDMD is computationally the most efficient, ResDMD and Hankel-DMD provide higher precision and robustness in spectral analysis at the expense of increased runtime, and EDMD-DL and ResKoopNet offer flexibility and accuracy through dictionary learning, with additional SGD iterations and optional pseudospectrum computation contributing to their computational burden.
@@
Empirical Perspective: In our experiments, without computing the pseudospectrum, the computational cost of ResDMD typically ranges from seconds to minutes. ResKoopNet, on the other hand, can require tens of minutes to several hours, depending on factors such as data dimensionality, the number of snapshots, hidden layer configurations, dictionary sizes, and training convergence criteria.
@@
Trade-off Between Cost and Accuracy: While ResKoopNet’s additional computational steps introduce higher costs, they enhance the accuracy and robustness of Koopman eigenpair estimation by allowing automatic dictionary learning and minimizing spurious spectral components. This trade-off makes ResKoopNet particularly valuable in applications where precision is critical. However, its computational demands render it less suitable for real-time or online Koopman model learning tasks.
Detailed Analysis of Computational Bottlenecks: The complexity $O(k(N_{K}^{2}m+N_{K}^{3}+|\theta|))$ arises from ResKoopNet’s iterative optimization, alternating between updating the neural network parameters $\theta$ and computing the Koopman matrix $\hat{K}(\theta)$. The dominant term $O(kN_{K}^{2}m)$ reflects the cost of constructing data matrices $\Psi_{X}(\theta)$ and $\Psi_{Y}(\theta)$ at each iteration, followed by gradient computation. For high-dimensional systems (large $m$) or complex architectures (large $|\theta|$), the number of iterations $k$ required for convergence increases, particularly if the loss $J(\theta)=\frac{1}{m}\|(\Psi_{Y}(\theta)-\Psi_{X}(\theta)K(\theta))V(\theta)\|_{F}^{2}$ is non-convex or ill-conditioned. The term $O(kN_{K}^{3})$ stems from solving $\hat{K}(\theta)=(G(\theta)+\sigma I)^{-1}A(\theta)$, where matrix inversion scales cubically with $N_{K}$. The optional pseudospectrum computation, with complexity $O(n_{z}N_{K}^{3})$, becomes significant for systems with continuous spectra, as $n_{z}$ grid points are scanned to evaluate residuals $\tau_{j}=\min_{\mathbf{v}_{i}}\overline{\operatorname{res}}(z_{j},\Psi\mathbf{v}_{i})$. Comparatively, ResDMD’s lower cost ($O(N_{K}^{2}m+n_{z}N_{K}^{3})$) avoids iteration, but sacrifices dictionary adaptability.
To mitigate these costs, we will consider some methods: (1) We consider Mini-Batch SGD since it uses a batch size $b\ll m$ that reduces per-iteration complexity to $O(N_{K}^{2}b+N_{K}^{3}+|\theta|)$, potentially lowering overall cost despite increased $k$, as validated by Barron space convergence properties. (2) We can use preconditioning to improve the condition number of $G(\theta)$ via diagonal scaling or incomplete Cholesky factorization could accelerate convergence, reducing $k$.
Theoretical Extensions: The $O(1/n)$ SGD rate assumes convexity, yet $J(\theta)$’s non-convexity suggests sublinear convergence. A bound like $\|\mathcal{K}-\mathcal{K}_{N_{K}}\|$ (Appendix A.4) could quantify accuracy, while determining minimal $N_{K}$ and $k$ for spectral completeness remains a future work. Extending ResKoopNet to stochastic systems (e.g., adding noise terms to spectral residual) or embedding physical constraints (e.g., unitary spectra for Hamiltonian systems) could enhance its scope, respectively, which also aligns with future work outlined in the Section 5.
A.7 Extra tests for ResDMD and ResKoopNet
A.7.1 ResDMD with less dictionary functions
While the original ResDMD implementation used 964 basis functions to capture the full spectrum of the nonlinear pendulum, our additional tests demonstrate that the unit circle spectrum can be reasonably approximated by ResDMD with fewer observables. Figure 7 shows the progressive improvement in spectral approximation as the dictionary size $N_{K}$ increases. With nearly 500 basis functions, ResDMD begins to adequately reconstruct the unit circle structure, though still requiring substantially more observables than the 300 basis functions needed by ResKoopNet.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) $N_{K}=252$ (b) $N_{K}=356$ (c) $N_{K}=464$ (d) $N_{K}=860$
Figure 7: Comparison across different dictionary sizes by ResDMD.
A.7.2 ResKoopNet with different hyperparameters
We test ResKoopNet on the dataset which has 90 initial points, as used in the Section 4.1. The Figure 8 shows that when tunning the number of hidden layers from 1 to 4 and the number of neurons in each layer from 250 to 350, the results change significantly. We can also tell that a suitable NN structure would be about 3 hidden layers with 300 neurons in each hidden layer. More neurons and hidden layers would be fore help, but cost more computational resource. This result also shows that our method is robust to hyperparameters tuning.
Refer to caption 1 hidden layer 250 neuron Refer to caption 1 hidden layer 275 neuron Refer to caption 1 hidden layer 300 neuron Refer to caption 1 hidden layer 325 neuron Refer to caption 1 hidden layer 350 neuron
Refer to caption 2 hidden layers 250 neuron Refer to caption 2 hidden layers 275 neuron Refer to caption 2 hidden layers 300 neuron Refer to caption 2 hidden layers 325 neuron Refer to caption 2 hidden layers 350 neuron
Refer to caption 3 hidden layers 250 neuron Refer to caption 3 hidden layers 275 neuron Refer to caption 3 hidden layers 300 neuron Refer to caption 3 hidden layers 325 neuron Refer to caption 3 hidden layers 350 neuron
Refer to caption 4 hidden layers 250 neuron Refer to caption 4 hidden layers 275 neuron Refer to caption 4 hidden layers 300 neuron Refer to caption 4 hidden layers 325 neuron Refer to caption 4 hidden layers 350 neuron
Figure 8: Comparisons of neural network architectures. Each row corresponds to the number of hidden layers (from 1 to 4), and each column shows the model with a different neuron count (250, 275, 300, 325, 350).
A.8 Hankel-DMD
A.8.1 Justification of using Hankel-DMD as comparison in all experiments
@@
As results, although its performance rivals ResKoopNet in the simple pendulum system by showing eigenvalues points near the unit circle and containing some polluted eigenvalues, which are close to the ground truth unit circle, we would like to emphasize that it capture the point spectrum and miss the full spectral information. When it comes to high-dimensional systems, it fails to capture key dynamics in higher-dimensional systems, as seen in the later experiment (Section 4.2 and Section 4.3).
A.8.2 Application in turbulence
Here we present the Koopman modes computed by Hankel-DMD for comparison with the ResKoopNet results. As shown in Figure 9, despite having small residuals, these modes fail to clearly capture the fundamental pressure field structure that was successfully identified by ResKoopNet’s first Koopman mode (see Figure 5). The discrepancy observed for Hankel-DMD [1] is due to its fundamentally different theoretical framework. While ResKoopNet (and other EDMD-based methods) use a Galerkin framework with convergence guarantees established via minimizing the spectral residual, the Hankel-DMD approach is based on Takens’ embedding theorem [36]. Therefore, the spectral residual metric, which is derived and justified within the Galerkin projection framework, does not fully extend to the Hankel-DMD setting. Even if Hankel-DMD yields small residual, this does not ensure that its modes accurately capture the underlying dynamics (such as the fundamental pressure field structure).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: The plots illustrate turbulence detection using the four Koopman modes computed by Hankel-DMD, which are ranked with their corresponding spectral residual from the smallest, similar as in ResKoopNet case.
A.9 Practical details for neural data analysis
A.9.1 Dataset details and experimental setup
@@
A.9.2 Task definition and rationale
In contrast to the competition’s prediction objective, our study focuses on the task of state partitioning of neural signals. While prediction remains feasible, we aim to demonstrate that state partitioning is sufficient to highlight the superiority of ResKoopNet over a series of other methods in uncovering the latent dynamics of the system. Specifically, in each experiment, a set of six video stimuli was repeatedly presented to each mouse, creating ideal conditions for defining brain states. The recording setup remained consistent for each mouse, ensuring that the neural activities could be interpreted as originating from the same dynamical system, with the primary variable being the input stimulus.
@@ A.9.2 Task definition and rationale
A.9.3 Dataset structure and dimensionality
@@
A.9.4 Implementations of ResKoopNet and other classical methods
We compare here four methods: the proposed ResKoopNet and three classical Koopman decomposition methods for high-dimensional systems: the Hankel-DMD, the EDMD with RBF basis, and the Kernel ResDMD. We applied them to the 5 datasets, although with slightly different implementations and different dimensions of approximated Koopman invariance subspace.
For ResKoopNet, we train the dictionaries with all the snapshots recorded in each mouse such that the total snapshot number is the product of the snapshot number in one trial and the number of all trials. This is to avoid overfitting with the small snapshot numbers within a trial. The high-dimensional data is first reduced to 300 dimensions with Singular Value Decomposition. The dimension of the Koopman subspace is chosen to be 50 (to be compared with Hankel DMD fairly), consisting of 25 trained bases and 25 pre-chosen ones (constant and the first-degree polynomials of the SVD-ed 24 dimensions). One can find the decomposed eigenfunctions in Figure 6A(top), with a marker of the ground truth state separations based on stimulus identity.
For Hankel-DMD, the Koopman eigenfunctions were approximated using the eigenvectors of the Hankel matrix. Specifically, the Hankel matrix was formed as in Equation 53 from Arbabi and Mezic [1], using all the observables from one trial of each mouse with a delay of 50. Consequently, the snapshot size became 249 times the observable number, and the resulting number of eigenfunctions was 50, each with a length of 50. The Hankel-DMD eigenfunctions for each trial of data are shown in Figure 6A (bottom), alongside the ground truth trial identities for comparison.
@@ A.9.4 Implementations of ResKoopNet and other classical methods
The Koopman eigenfunctions from both ResKoopNet and other methods represent dynamical features corresponding to one of the six video stimuli. To evaluate how well the eigenfunctions capture the latent dynamics, we assess the similarity of the features for trials with the same stimulus and their dissimilarity from those corresponding to different stimuli. Effectively, this makes the problem a clustering task, where the separability of the Koopman eigenfunctions reflects how well they capture the key dynamic components related to the stimuli.
A.10 Choice justification of dictionary sizes
@@
    For ResKoopNet, we selected 25 trained basis functions and 24 first-order monomial basis functions as the dictionary for the 24 reduced observables. This choice ensures the dictionary size is the same as Hankel DMD, making it a fair comparison between the performance of these two methods. The last two methods uses higher dictionary sizes but still remain ineffective in capturing temporal dynamics, thus not affecting the fairness of choosing 50 dictionaries here. The size of the trained dictionary was set to be at least equal to the original observable size to ensure a sufficiently rich dictionary for the Koopman invariant subspace.
@@
    For Hankel DMD, the number of delays (as dictionary size/number of eigenfunctions) is first constrained by the temporal sample size (i.e., snapshot size) because it cannot exceed the maximum snapshot size. Therefore, it is impossible to choose the same dictionary size as the ResKoopNet example. Choosing the delay too small will result in an insufficient dictionary size to span the Koopman invariant subspace, and too large will reduce the actual snapshot size to estimate the covariance matrices in the estimation of the Koopman matrix. Therefore, we chose a compromise delay number of 50 that satisfies both needs.
@@
    For RBF basis, in principle, we can use the same dictionary size. However, our previous experience with a similar dataset and the results of using the RBF basis for the EDMD method all suggest that the performance will be better with more dictionary functions. Therefore, we chose 1000 RBF basis and the original 300 first-order monomial basis as a better condition compared to the same dictionary size with ResKoopNet.
@@
    For Kernel ResDMD, the dictionary size is theoretically determined to be the number of snapshots. Therefore, we cannot make the dictionary size consistent with the ResKoopNet example.
@@
A.10.1 Visualization and clustering performance